# A Comprehensive Survey on Uncertainty Quantification for Deep Learning

### v4

@@ 6.1. Out-of-distribution detection
Given a training data distribution $p(x)$, the OOD data are those samples that are either unlikely under the training data distribution or outside the support of $p(x)$. Accurate detection of OOD samples is of paramount importance for the safety-critical application, e.g., autonomous driving (Shafaei et al., 2018), medical image analysis (Ghoshal et al., 2021) Since OOD samples lie further away from the training samples, the trained model may not generalize well and produces diverse and unstable predictions for those samples. The model parameters or architecture are not suitable for OOD samples. Thus OOD samples are expected to have higher model uncertainty. The primary uncertainty for OOD data is concerned with model uncertainty, because the model trained with in-distribution may not generalize well to other domains. It is important to make the model aware of the sample density when quantifying the uncertainty.
Existing approaches: From the above analysis we are clear that model uncertainty is more important for OOD detection. Thus existing approaches leveraging the model uncertainty framework are much more popular. For example, drop-out-based BNN approaches have been applied to OOD detection and improved the performance through the use of the randomized embeddings induced by the intermediate layers of a dropout BNN (Nguyen et al., 2022a); Deep ensembles are simple but perform well on OOD detection (Lakshminarayanan et al., 2017; Ovadia et al., 2019). Recent advances attempt to develop distance-aware DNN for more accurate OOD detection by imposing constraints on the feature extracting process (Liu et al., 2020; van Amersfoort et al., 2021). Recent developments in evidential deep learning framework have also demonstrated its capability on OOD detection in many benchmark datasets because of the explicit distinction between two types of uncertainty in the framework (Malinin and Gales, 2018; Charpentier et al., 2020; Sensoy et al., 2018; Bao et al., 2021).
@@ 6.3. Deep Reinforcement Learning
The purpose of deep reinforcement learning (DRL) is to train an agent interacting with environment to maximize its total rewards (Sutton and Barto, 2018). DRL can be regarded as learning via Markov Decision Process, defined by the tuple $\{S,A,R,P\}$, where $S$ is the set of states (environment condition), $A$ is the set of actions (agent), $R$ is the function mapping state-action pairs to rewards, and $P$ is the transition probability (the probability of next state after performing actions on current state). The goal of DRL is to learn the policy $\pi$ (a function mapping given state to an action) that maximize the sum of discounted future rewards $J(\pi)=\mathbb{E}[\sum_{i}\gamma^{i}R(s_{i},a_{i})]$, where $\gamma$ is the discount factor on the future reward (meaning rewards in more future are less important) (Sutton and Barto, 2018). In Deep Q-learning, the DNN models are utilized to learn the value functions (expected rewards for a state-action pair) for each action given the input state.
@@
Due to the complex agent and environment conditions and limited training states, there also exist two types of uncertainty sources: data and model uncertainty. Data uncertainty can arise from the intrinsic randomness in the interactions between the agent and environment, which causes randomness in the reward $R$, the transition functions $P$ and the randomness in next state value distribution. To characterize the data uncertainty arising from those sources, the distributional RL (Bellemare et al., 2017) take a probabilistic perspective on learning the rewards functions instead of approximating the expectation of the value. Thus the approach can implement risk-aware behavior on the agent. Similar approach is proposed to quantify the data uncertainty in DRL aiming for curiosity-based learning in face of unpredictable transitions (Mavor-Parker et al., 2022). The following work (Dabney et al., 2018) extend the parametric distribution to non-parametric prediction interval methods to quantify the data uncertainty and avoid the explicit parametric format. The approach corresponds to the literature we discuss in section 5.1. On the other hand, given the limited training state space, there exists high model uncertainty. The DNN model may not learn the optimum policy function and miss the unexplored state spaces, which potentially give higher rewards. This means DRL model face a trade-off between exploitation and exploration. Exploration means utilizing what the model has learned and choose the best policy to maximize future rewards. Exploration refers to choose the unexplored state to learn more possible high state-action pairs. The challenge of effective exploration is connected to model uncertainty. The higher model uncertainty means the model are not learnt well on the given state and require more exploration on that sample. For example, deep ensemble Q-network (Chen et al., 2021c) is proposed to inject model uncertainty into Q learning for more efficient exploration sampling. To reduce the computational overhead, Dropout Q-functions (Hiraoka et al., 2021) method is proposed to use MC-dropout for model uncertainty quantification. The following work (Osband et al., 2018) demonstrates the previous ensemble and dropout methods may produce poor approximation to the model uncertainty in case where state density does not correlate with the true uncertainty. To overcome the shortcoming, they suggest adding a random prior to the ensemble DQNs.
@@
In summary, both data and model uncertainty are crucial aspects of DRL for more efficient policy learning and exploration. Many works inspired from DNN uncertainty quantification have been proposed to improve the performance of DRL.
@@ 7.1. Combine UQ with DNN explainability
The explanation for DNN model predictions has been increasingly crucial because it provides tools for understanding and diagnosing the model’s prediction. Recently, many explainability methods, termed explainers (Vu and Thai, 2020), have been introduced in the category of local feature attribution methods. That is, methods that return a real-valued score for each feature of a given data sample, representing the feature’s relative importance for the sample prediction. These explanations are local in that each data sample may have a different explanation. Using local feature attribution methods, therefore, helps users better understand nonlinear and complex black-box models. Both uncertainty quantification and explanation are important for a robust, trustworthy AI model. Current methodologies consider two directions separately, we consider it could enable a more trustworthy AI system if combining them. Though many methodologies have been proposed for more precise uncertainty quantification, very few techiniques attempt to explain why the uncertainty exists in the predictions.
There are two possible directions to combine the power of explanations and uncertainty quantification: First, existing explanation methods could be potentially improved after considering the prediction uncertainty since those uncertain samples’ explanations may not be trustworthy and can be omitted. Second, from another perspective, after obtaining the uncertainty quantification, we can leverage the existing post-hoc explanation methods for understanding the reason that the model is uncertain. For example, it is intriguging to ask the question why the prediction is uncertainty and which or which set of input features are uncertain, or due to which layer of the model is imperfect.
@@ 7.2. Structured Uncertainty Quantification
Structured data are samples that are interdependent with each other, violating the common i.i.d assumption (BakIr et al., 2007). Examples are imaging data, spatiotemporal data, and graphs. Deep learning has been widely used to model structured data, but the uncertainty of its prediction is not often quantified. Here, we list future research directions for the three different types of structured data.
@@ 7.2.2. Spatiotemporal data
Spatiotemporal data are special due to the violation of the common assumption that samples follow an identical and independent distribution  (Shekhar et al., 2015; Jiang, 2018; Jiang et al., 2022b). Uncertainty quantification of spatiotemporal deep learning poses several unique challenges. First, the analysis of spatiotemporal data requires the co-registration of different maps (e.g., points, lines, polygons, geo-raster) into the same spatial reference system. The process is subject to registration uncertainty due to GPS errors or annotation mistakes in map generation  (Jiang et al., 2022a, 2021). Such registration uncertainty causes troubles when training deep neural networks  (He et al., 2022). Second, implicit dependency structures exist in continuous space and time (e.g., spatial and temporal autocorrelation, and temporal dynamics). Thus, the uncertainty quantification process should be aware of such a dependency structure. Third, spatiotemporal non-stationary requires characterizing uncertainty due to out-of-distribution samples (Shekhar et al., 2015; Jiang, 2018). In addition, a different level of uncertainty exists based on the nearby training sample density. Traditionally, the Gaussian process has been widely used to quantify spatial uncertainty. However, for deep neural network models, new techniques are needed that consider sample density both in the non-spatial feature space and in the geographic space.
@@ 7.3. UQ for physics-aware DNN models
Many scientific applications can be described with a physics principle, e.g., the conservation law for diffusion process in the form of a partial differential equation (PDE). Deep learning shows a great capability of extracting complex patterns from massive data in scientific discoveries. However, black-box deep neural networks do not incorporate the fundamental physics principle of the system (Kashinath et al., 2021). To address the challenge, physics-informed neural networks (PINN) (Krishnapriyan et al., 2021; Yang and Perdikaris, 2019; Li et al., 2020; Karniadakis et al., 2021) attempt to incorporate physical principles and domain knowledge into deep learning to make physically consistent predictions. For example, considering a spatio-temporal PDE in the form of
@@
For many physical systems, the underlying physical principle generating the samples may be non-deterministic or unknown. The system is inherently chaotic and stochastic. For example, in weather forecasting and climate prediction problem, the state of the system is sensitive to the initial conditions, in which a small perturbation in the nonlinear system can result in a large difference in the later state (a.k.a. butterfly effect) (Kashinath et al., 2021). Even with a deterministic initial condition, the unknown physics equation (e.g., PDE) that drives the system can be stochastic, which is described by a stochastic differential equation. Uncertainty quantification is of vital importance to improve the prediction’s reliability, especially under distribution shifts. To summarize, the uncertainty of physical system modeling can come from the following source. First, the initial and boundary condition of the physical system is non-deterministic, and the system may be chaotic (Wang et al., 2019b). Second, the inherent physical principle may not be perfectly known or the parameter of the governing equation may be stochastic, i.e., in an imperfect physical system, the conservation law of heat may be violated in a non-closed system (Yang and Perdikaris, 2019; Daw et al., 2021).
These cases indicate that for the stochastic physical system, there exists inherent data uncertainty. Thus, a natural way to model distributions and incorporate stochasticity and uncertainty into the neural network is through probabilistic models. Because of the unique characteristic of physical data, there exist several challenges for quantifying the uncertainty in PINN. First, the system requires considering the physical principle and its uncertainty simultaneously. The incorporation of physical constraints can significantly reduce the data and model uncertainty. Second, there are various sources of uncertainty, which may arise from the inherent physical system randomness, the measurement error, and the limited knowledge of the physical governing equation.
@@ 8. Conclusion
Acknowledgements.
This material is based upon work supported by the National Science Foundation (NSF) under Grant No. IIS-2147908, IIS-2207072, CNS-1951974, OAC-2152085, and the National Oceanic and Atmospheric Administration grant NA19NES4320002 (CISESS) at the University of Maryland.

### v5

@@ 6.1. Out-of-distribution detection
Given a training data distribution $p(x)$, the OOD data are those samples that are either unlikely under the training data distribution or outside the support of $p(x)$. Accurate detection of OOD samples is of paramount importance for safety-critical applications, e.g., autonomous driving (Shafaei et al., 2018), medical image analysis (Ghoshal et al., 2021) Since OOD samples lie further away from the training samples, the trained model may not generalize well and produces diverse and unstable predictions for those samples. The model parameters or architecture are not suitable for OOD samples. Thus OOD samples are expected to have higher model uncertainty. The primary uncertainty for OOD data is concerned with model uncertainty, because the model trained with in-distribution may not generalize well to other domains. It is important to make the model aware of the sample density when quantifying the uncertainty.
Existing approaches: From the above analysis we are clear that model uncertainty is more important for OOD detection. Thus existing approaches leveraging the model uncertainty framework are much more popular. For example, drop-out-based BNN approaches have been applied to OOD detection and improved the performance through the use of the randomized embeddings induced by the intermediate layers of a dropout BNN (Nguyen et al., 2022a) and node-based BNN (Trinh et al., 2022); Deep ensembles are simple but perform well on OOD detection (Lakshminarayanan et al., 2017; Ovadia et al., 2019; Zaidi et al., 2021). Recent advances attempt to develop distance-aware DNN for more accurate OOD detection by imposing constraints on the feature extracting process (Liu et al., 2020; van Amersfoort et al., 2021). Recent developments in evidential deep learning framework have also demonstrated its capability on OOD detection in many benchmark datasets because of the explicit distinction between two types of uncertainty in the framework (Malinin and Gales, 2018; Charpentier et al., 2020; Sensoy et al., 2018; Bao et al., 2021).
@@ 6.3. Deep Reinforcement Learning
The purpose of deep reinforcement learning (DRL) is to train an agent interacting with the environment to maximize its total rewards (Sutton and Barto, 2018). DRL can be regarded as learning via the Markov Decision Process, defined by the tuple $\{S,A,R,P\}$, where $S$ is the set of states (environment condition), $A$ is the set of actions (agent), $R$ is the function mapping state-action pairs to rewards, and $P$ is the transition probability (the probability of next state after performing actions on current state). The goal of DRL is to learn the policy $\pi$ (a function mapping given state to an action) that maximizes the sum of discounted future rewards $J(\pi)=\mathbb{E}[\sum_{i}\gamma^{i}R(s_{i},a_{i})]$, where $\gamma$ is the discount factor on the future reward (meaning rewards in more future are less important) (Sutton and Barto, 2018). In Deep Q-learning, the DNN models are utilized to learn the value functions (expected rewards for a state-action pair) for each action given the input state.
@@
Due to the complex agent and environment conditions and limited training states, there also exist two types of uncertainty sources: data and model uncertainty. Data uncertainty can arise from the intrinsic randomness in the interactions between the agent and environment, which causes randomness in the reward $R$, the transition functions $P$, and the randomness in the next state value distribution. To characterize the data uncertainty arising from those sources, the distributional RL (Bellemare et al., 2017) takes a probabilistic perspective on learning the rewards functions instead of approximating the expectation of the value. Thus the approach can implement risk-aware behavior on the agent. A similar approach is proposed to quantify the data uncertainty in DRL aiming for curiosity-based learning in the face of unpredictable transitions (Mavor-Parker et al., 2022). The following work (Dabney et al., 2018) extends the parametric distribution to non-parametric prediction interval methods to quantify the data uncertainty and avoid the explicit parametric format. The approach corresponds to the literature we discuss in section 5.1. On the other hand, given the limited training state space, there exists high model uncertainty. The DNN model may not learn the optimum policy function and miss the unexplored state spaces, which potentially give higher rewards. This means the DRL model faces a trade-off between exploitation and exploration. Exploration means utilizing what the model has learned and choosing the best policy to maximize future rewards. Exploration refers to choosing the unexplored state to learn more about possible high state-action pairs. The challenge of effective exploration is connected to model uncertainty. The higher model uncertainty means the model is not learned well in the given state and requires more exploration of that sample. For example, deep ensemble Q-network (Chen et al., 2021c) is proposed to inject model uncertainty into Q learning for more efficient exploration sampling. To reduce the computational overhead, the Dropout Q-functions (Hiraoka et al., 2021) method is proposed to use MC-dropout for model uncertainty quantification. The following work (Osband et al., 2018) demonstrates the previous ensemble and dropout methods may produce a poor approximation to the model uncertainty in cases where state density does not correlate with the true uncertainty. To overcome the shortcoming, they suggest adding a random prior to the ensemble DQNs.
@@
In summary, both data and model uncertainty are crucial aspects of DRL for more efficient policy learning and exploration. Many works inspired by DNN uncertainty quantification have been proposed to improve the performance of DRL.
@@ 7.1. Combine UQ with DNN explainability
The explanation for DNN model predictions has been increasingly crucial because it provides tools for understanding and diagnosing the model’s prediction. Recently, many explainability methods, termed explainers (Vu and Thai, 2020), have been introduced in the category of local feature attribution methods. That is, methods that return a real-valued score for each feature of a given data sample, representing the feature’s relative importance for the sample prediction. These explanations are local in that each data sample may have a different explanation. Using local feature attribution methods, therefore, helps users better understand nonlinear and complex black-box models. Both uncertainty quantification and explanation are important for a robust, trustworthy AI model. Current methodologies consider two directions separately, we consider it could enable a more trustworthy AI system if combining them. Though many methodologies have been proposed for more precise uncertainty quantification, very few techniques attempt to explain why uncertainty exists in the predictions.
There are two possible directions to combine the power of explanations and uncertainty quantification: First, existing explanation methods could be potentially improved after considering the prediction uncertainty since those uncertain samples’ explanations may not be trustworthy and can be omitted (Zhao et al., 2021; Zhang et al., 2022). Second, from another perspective, after obtaining the uncertainty quantification, we can leverage the existing post hoc explanation methods to understand the reason that the model is uncertain (Slack et al., 2021). For example, it is intriguing to ask the question of why the prediction is uncertain and which set of input features are uncertain, or due to which layer of the model is imperfect.
@@ 7.2. Structured Uncertainty Quantification
Structured data are samples that are interdependent with each other, violating the common i.i.d assumption (BakIr et al., 2007; lyu2022study). Examples are imaging data, spatiotemporal data, and graphs. Deep learning has been widely used to model structured data, but the uncertainty of its prediction is not often quantified. Here, we list future research directions for the three different types of structured data.
@@ 7.2.2. Spatiotemporal data
Spatiotemporal data are special due to the violation of the common assumption that samples follow an identical and independent distribution  (Shekhar et al., 2015; Jiang, 2018; Jiang et al., 2022b; deng2024incremental). Uncertainty quantification of spatiotemporal deep learning poses several unique challenges. First, the analysis of spatiotemporal data requires the co-registration of different maps (e.g., points, lines, polygons, geo-raster) into the same spatial reference system. The process is subject to registration uncertainty due to GPS errors or annotation mistakes in map generation  (Jiang et al., 2022a, 2021). Such registration uncertainty causes troubles when training deep neural networks  (He et al., 2022). Second, implicit dependency structures exist in continuous space and time (e.g., spatial and temporal autocorrelation, and temporal dynamics). Thus, the uncertainty quantification process should be aware of such a dependency structure. Third, spatiotemporal non-stationary requires characterizing uncertainty due to out-of-distribution samples (Shekhar et al., 2015; Jiang, 2018). In addition, a different level of uncertainty exists based on the nearby training sample density. Traditionally, the Gaussian process has been widely used to quantify spatial uncertainty. However, for deep neural network models, new techniques are needed that consider sample density both in the non-spatial feature space and in the geographic space.
@@ 7.3. UQ for physics-aware DNN models
Many scientific applications can be described with a physics principle, e.g., the conservation law for diffusion process in the form of a partial differential equation (PDE). Deep learning shows a great capability of extracting complex patterns from massive data in scientific discoveries. However, black-box deep neural networks do not incorporate the fundamental physics principle of the system (Kashinath et al., 2021). To address the challenge, physics-informed neural networks (PINN) (Krishnapriyan et al., 2021; Yang and Perdikaris, 2019; Li et al., 2020; Karniadakis et al., 2021) attempt to incorporate physical principles and domain knowledge into deep learning to make physically consistent predictions. For example, considering a spatiotemporal PDE in the form of
@@
For many physical systems, the underlying physical principle generating the samples may be non-deterministic or unknown. The system is inherently chaotic and stochastic. For example, in weather forecasting and climate prediction problems, the state of the system is sensitive to the initial conditions, in which a small perturbation in the nonlinear system can result in a large difference in the later state (a.k.a. butterfly effect) (Kashinath et al., 2021). Even with a deterministic initial condition, the unknown physics equation (e.g., PDE) that drives the system can be stochastic, which is described by a stochastic differential equation. Uncertainty quantification is of vital importance to improve the prediction’s reliability, especially under distribution shifts. To summarize, the uncertainty of physical system modeling can come from the following source. First, the initial and boundary conditions of the physical system are non-deterministic, and the system may be chaotic (Wang et al., 2019b). Second, the inherent physical principle may not be perfectly known or the parameter of the governing equation may be stochastic, i.e., in an imperfect physical system, the conservation law of heat may be violated in a non-closed system (Yang and Perdikaris, 2019; Daw et al., 2021).
These cases indicate that for the stochastic physical system, there exists inherent data uncertainty. Thus, a natural way to model distributions and incorporate stochasticity and uncertainty into the neural network is through probabilistic models. Because of the unique characteristics of physical data, there exist several challenges for quantifying the uncertainty in PINN. First, the system requires considering the physical principle and its uncertainty simultaneously. The incorporation of physical constraints can significantly reduce the data and model uncertainty. Second, there are various sources of uncertainty, which may arise from the inherent physical system randomness, the measurement error, and the limited knowledge of the physical governing equation.
@@
7.4. UQ for Large Language Models
In recent years, large language models (LLMs) have demonstrated remarkable capabilities in generating coherent and contextually relevant text across various applications, from natural language processing to decision-making tasks. These models, such as OpenAI’s GPT-4 (Achiam et al., 2023), have shown proficiency in tasks ranging from text summarization and translation to more complex problem-solving and creative writing. However, despite their impressive performance, a critical challenge remains: how to ensure accurate uncertainty quantification to enhance the trustworthiness of these models? Existing research in LLMs uncertainty quantification mainly falls into two main categories: uncertainty estimation, and confidence calibration (Geng et al., 2023). Uncertainty estimation methods for LLMs includes logit-based methods (Duan et al., 2023), model internal state-based methods (Ren et al., 2022), and consistency-based estimation (Manakul et al., 2023b). Logit-based methods use the logits in the model output layer to compute the class entropy of output tokens. Internal state-based methods use the variability of sample embedding from model intermediate layers to quantify uncertainty. The embedding that is far from the majority of others indicates high uncertainty. Consistency-based methods measure the consistency between different model outputs from a sampling process and use the consistency as the uncertainty score (Kotelanski et al., 2023; Manakul et al., 2023b; Lin et al., 2023). For uncertainty estimation in LLMs, it is important to consider the semantic equivalence of the generated answers as different model outputs may look different in the generated tokens but similar in semantics. Graph-based (Lin et al., 2023) and semantic methods (Kuhn et al., 2023) propose alternative strategies for confidence estimation based on the semantic pattern, using a semantics similarity graph of different model outputs for uncertainty quantification. For confidence calibration, the goal is to adjust the model so that the quantified uncertainty from the above methods is aligned with prediction accuracy (Mielke et al., 2022; Liu et al., 2024a). However, the field is still in its infancy. In future research, more work needs to be done to incorporate the natural ambiguity of language. For example, the same token sequences may have different semantic meanings in different contexts. As discussed in (Geng et al., 2023), benchmarking datasets and UQ methods for multi-modal foundation models are missing.